OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the circumference of a circle, C = 2 piR?

OpenStudy (anonymous):

C/2r= pi 2pi/c=r C/2Pi=r C/rpi=2

OpenStudy (shubhamsrg):

hmmn... tell me this..if 8 =2 * 4 which one is equivalent ? a) 8/2 = 4 b) 8/4 =2 c) 2*4 = 8 d) none of these ?????????????????????

OpenStudy (anonymous):

That doesnt have anything to do with my question

OpenStudy (anonymous):

b

OpenStudy (shubhamsrg):

really?? o.O

OpenStudy (anonymous):

C=2.PI/R

OpenStudy (anonymous):

Raphael thats the circumfrence

OpenStudy (anonymous):

What shubhamsrg is trying to do is show you how to find the answer, if you substitute numbers into the equation you will be able to understand why b is a false statement

OpenStudy (shubhamsrg):

hmmn.. @bbsmilieface ! finally!! someone's there!! :'( :P

OpenStudy (anonymous):

ooo ok thnk you i didnt understand what he ment

OpenStudy (anonymous):

sorry @shubhamsrg lol i didnt understand it at first

OpenStudy (shubhamsrg):

really? whats your age buddy?

OpenStudy (anonymous):

14

OpenStudy (shubhamsrg):

hmmn..i wont comment on this,,or infact..i shouldnt comment on ths.. :|

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

ok so which one is it lol still havnt gotten that the number thing doesnt help me at all

OpenStudy (anonymous):

Nvr mnd i got